You initially purchase 150 shares at $15 per share, you decided to reinvest your distributions and end with 175 shares worth $18.50 a share. What is your total return?

  1. Why would you choose to reinvest distributions instead of taking them out?
Answer #1

Value of investment = Market price of Share * No. of share

Invested amount = $15 * 150 share = $2,250

After reinvesting the distributed amount...

The current value of the investment = $18.5 * 175 Share = $3,237.5

Return on investment in % = (Current value of investment - invested amount ) * 100 / invested amount

Return on Investment in % = (3237.5 - 2250) * 100 / 2250

= 43.89%

  

Why would you choose to reinvest the distributed amount?

We have two option after receiving a distributed amount

1.) To reinvest that amount in the same stock or

2.) To reinvest that amount somewhere else or do use for consumption.

When we are expecting higher returns/utility from reinvesting in the same stock/company compared to another option we will reinvest distributed amounts in the same stock/company.